OpenStudy (mimi_mik):

Help! 8th grade algebra homework. n+5(n-1)=7

OpenStudy (candystarlover2):

https://scribblar.com/mkyvzn9j I'll help :) just click the link

OpenStudy (candystarlover2):

:)

OpenStudy (candystarlover2):

the link I added has the answer

OpenStudy (turbostudier):

Here is how to do the problem: n+ 5n-5=7

OpenStudy (turbostudier):

Do you know how i did that?

OpenStudy (turbostudier):

@mimi_mik

OpenStudy (turbostudier):

i rewrote the problem, except i solved for what was within the parenthesis. since the 5 was on the outside, i multiplied the 5 by everything inside the parenthesis

OpenStudy (turbostudier):

so you get n+5n-5=7

OpenStudy (turbostudier):

but we still have to solve

OpenStudy (turbostudier):

so then you add like terms: n+5n=6n then you bring the 5 to the other side to join the 7, and you use inverse operations. So then its 6n=7+5 6n=12 n=2

OpenStudy (turbostudier):

And that is how you solve your problem!! ;D

OpenStudy (turbostudier):

So make sure you work through tour problems like that. Its like untying a really big knot. you analyze it, then try pulling different threads to see if it will break loose!!!;D (Good analogy)And good luck!

OpenStudy (mimi_mik):

@turbostudier thank you!